OpenStudy (callisto):

Express t in terms of x: t = \frac{x^2-t+1}{x+t+1}

ganeshie8 (ganeshie8):

I'd rearrange and use quadratic formula

OpenStudy (callisto):

Attempt: t(x+t+1) = x^2 - t + 1 xt + t^2 + 2t = x^2 + 1 t ( x + t + 2) = x^2 + 1 And then I got stuck...... Any help would be greatly appreciated!

OpenStudy (callisto):

I cannot use the quadratic formula, since it is not taught at that level...

ganeshie8 (ganeshie8):

Its a quadratic in t : t^2 + (x+2)t -x^2 -1 = 0

OpenStudy (callisto):

So? I still cannot use the quadratic formula, as the ones who attempt this question have not learnt the quadratic formula yet!

sam (.sam.):

\[t=\frac{x^2-t+1}{x+t+1} \\ \\ t(x+t+1)=x^2-t+1 \\ \\ tx+t^2+t=x^2-t+1 \\ \\ t^2+xt+2t=x^2+1 \\ \\ t^2+(x+2)t=x^2+1 \] Completing the square method? \[t^2+(x+2)t+\frac{(x+2)^2}{4}=x^2+1+\frac{(x+2)^2}{4} \\ \\ (t+\frac{x+2}{2})^2=x^2+1+\frac{(x+2)^2}{4} \\ \\ t=\frac{-x-2}{2}\pm \sqrt{x^2+1+\frac{(x+2)^2}{4}}\]

OpenStudy (callisto):

t = (x^2-t+1) / (x+t+1) t(x+t+1) = x^2-t+1 xt + t^2 + t = x^2 - t + 1 t^2 + (x+2)t = x^2 + 1 t^2 + (x+2)t + [(x+2)/2]^2 = x^2 + 1 + [(x+2)/2]^2 {t + [(x+2)/2]}^2 = 5x^2/4 + x + 3/2 t + [(x+2)/2] =+/- sqrt(5x^2/4 + x + 3/2) t = -x/2 - 1 +/- sqrt(5x^2/4 + x + 3/2) That's the best I can do with completing square, but it does look ugly.

sam (.sam.):

Yep

OpenStudy (callisto):

I don't suppose they know that. That's too much for them... I didn't learn this at that level as well...

OpenStudy (kainui):

Are you sure you wrote the problem down correctly? Does seem unusually hard for the level it's given to.

OpenStudy (kainui):

Like for instance what if it wasn't this: \[t =\frac{x^2-t+1}{x+t+1}\] but instead two parenthesis and a square was left off as an accident: \[t =\frac{x^2-(t+1)^2}{x+t+1}\] then it becomes: \[t= \frac{(x+t+1)(x-t-1)}{x+t+1} = x-t-1\] now it's super easy: \[t=\frac{x-1}{2}\] OK that's probably not what happened though lol.

ganeshie8 (ganeshie8):

Yeah, I have the same feeling, that top must be a difference of squares to make the solution simple

OpenStudy (callisto):

I will have to ask my math teacher to find it out tomorrow. Thanks everyone for your time!

ganeshie8 (ganeshie8):

Do let us know if your teacher has any interesting way to handle this. Seems unlikely though

OpenStudy (callisto):

Update: My teacher said it is too difficult for them as it is quadratic, therefore she cancelled the question. Thanks everyone for your help! :)